# Sliding Speed Influences Electrovibration-Induced Finger Friction Dynamics on Touchscreens

### v1

@@ 5 Conclusions
Our results show that finger stiffness increases more than moving mass as sliding speed increases, leading to a rise in the cutoff frequency of the electrovibration-induced friction response. We also observed that low and moderate applied forces do not influence skin mechanics or finger electrovibration response. Building on these insights, we developed a practical, speed-dependent electrovibration-induced friction model that dynamically adjusts the input voltage to achieve the desired output friction. This model enables researchers to enhance the realism of texture rendering algorithms and ensure consistent haptic feedback for users by minimizing the discrepancy between the input intended texture spectrum and the output electrovibration-induced friction spectrum.
@@ A.2 Generated friction force through amplitude modulated input signal in electrovibration
For amplitude modulation, we multiply the above equation by a high-frequency carrier wave, $f_{c}$, usually beyond 5 kHz. The DC shift, $\min|V_{m}(t)|$, causes the amplitude modulation to be a double-sideband full carrier. The peak amplitude of $V_{m}(t)$ and carrier is 1V, resulting in a modulation index of 1. The modulation index should be $<=1$ to prevent signal distortion due to over-modulation. We chose the modulation index as unity to ensure the input message signal’s power is higher. The carrier signal does not contain useful information for texture perception since the electrovibration friction intensity depends on the input voltage signal’s, $V_{i}(t)$, power.
@@
Since the finger acts as a low-pass filter with a cut-off at 1 kHz, the higher-frequency components are attenuated:
[…]
@@
Appendix B Appendix B
B.1 Setup response
@@
$\frac{v_{snd}(s)}{F_{snd}(s)}=H_{snd}(s)=\frac{sK_{snd}\omega_{n,snd}^{2}}{s^{2}+2\zeta_{snd}\omega_{n,snd}s+\omega_{n,snd}^{2}},$ (22)
@@
B.2 Removing setup response

### v2

@@ 5 Conclusions
Our results demonstrate that, within the sliding range of 20–100 mm/s, finger stiffness increases by 4.04 N/m for each unit increase in speed (relative change of 1.11%), which exceeds the corresponding increase in moving mass, measured at $3.23\times 10^{-5}$ kg (0.82%). This disparity results in a 13.8 Hz (2.47%) rise in the cutoff frequency of the electrovibration-induced finger friction response. Furthermore, we found that low to moderate applied forces (approximately 0.1 N to 0.6 N) had no statistically significant effect on either the skin’s mechanical properties or the parameters of the electrovibration-induced finger friction response. Based on these findings, we developed a practical, speed-dependent model that dynamically adjusts the input voltage to achieve the desired output friction. This model enables researchers to enhance the realism of texture rendering algorithms by minimizing discrepancies between the target texture spectrum and the resulting electrovibration-induced finger friction spectrum, thereby supporting more consistent and perceptually accurate haptic feedback.
@@ 8 Declaration of competing interest
9 Declaration of generative AI and AI-assisted technologies in the writing process
The authors used OpenAI’s ChatGPT and Grammarly Inc.’s Grammarly tool to enhance the language and read- ability of this work. After using these tools, the authors carefully reviewed and edited the content as needed. They take full responsibility for the final publication.
@@ A.2 Generated friction force through amplitude modulated input signal in electrovibration
For amplitude modulation, we multiply the above equation by a high-frequency carrier wave, $f_{c}$, usually beyond 5 kHz. The DC shift, $\min|V_{m}(t)|$, causes the amplitude modulation to be a double-sideband full carrier. The peak amplitude of $V_{m}(t)$ and carrier is 1V, resulting in a modulation index of 1. The modulation index should be <=1<=1 to prevent signal distortion due to over-modulation. We chose the modulation index as unity to ensure the input message signal’s power is higher. The carrier signal does not contain useful information for texture perception since the electrovibration friction intensity depends on the input voltage signal’s, $V_{i}(t)$, power.
@@
Since the finger acts as a low-pass filter with a cutoff at 1 kHz, the higher-frequency components are attenuated:
$F_{e}=\frac{k}{2}(V_{m}\cos(2\pi f_{m}t)+\min|V_{m}(t)|+\cancel{V_{m}(\cos(2\pi f_{c}+f_{m}t)+\cos(2\pi f_{c}-f_{m}t))}$ (21)
$+\cancel{\min|V_{m}(t)|\cos(2\pi f_{c}t))}$
@@
Appendix B Setup response
@@
$\frac{v_{snd}(s)}{F_{snd}(s)}=H_{snd}(s)=\frac{sK_{snd}\omega_{n,snd}^{2}}{s^{2}+2\zeta_{snd}\omega_{n,snd}s+\omega_{n,snd}^{2}}$ (23)
@@
B.1 Removing setup response
@@
Appendix C Exploration condition check
During the experiment, we ensured that participants adhered to the specified exploration conditions in each trial. To verify this, we recorded both the applied force and the scanning speed of the participant’s finger throughout each trial. Applied force data were sampled at 20 kHz, and scanning speed at 60 Hz, resulting in 200,000 and 600 samples over a 10-second trial. Threshold ranges were defined as $\pm 10\%$ for applied force and $\pm 25\%$ for sliding speed, relative to target values.
For each trial, we first counted the number of samples that exceeded the $\pm 25\%$ threshold for each parameter. Additionally, we computed the mean applied force and sliding speed and verified whether these values remained within $\pm 10\%$ of the corresponding targets. A trial was deemed invalid and repeated if either of the following conditions was met: 1) the number of out-of-threshold samples exceeded the equivalent of one second of data—i.e., more than 20,000 samples for applied force or more than 60 samples for scanning speed—or 2) the mean value fell outside the $\pm 10\%$ range. Figures 12 and 13 present representative data for applied force and sliding speed from a single trial, illustrating the verification process.
Figure 12: Illustration of force applied by a participant in a trial for the 0.2 N target setting.
Figure 13: An illustration of the finger scanning speed recorded during a trial conducted with a 40 mm/s target condition.
Appendix D Finger exploration trajectory
Participants used the index finger of their dominant hand to explore the touchscreen in lateral and medial directions. Figure 14 presents a heatmap of finger position data, sampled at 60 Hz across all participants and trials. The infrared position sensor covered an active area of 135 × 158 mm, with the 35 × 70 mm touchscreen centered within this area.
Figure 14: Heatmap of the finger exploration trajectory of all participants across all trials. The cyan box indicates the dimensions of the touchscreen.
Appendix E Percentage contribution of applied force and scanning speed on GLMM model outputs
Consider, for example, the equation for gain, $K$, modeled for fixed effects of applied force, $F_{n}$, and scanning speed, $v$, and corresponding coefficients, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ for applied force, scanning speed, and their interaction, respectively.
For the percentage contribution calculation, we compute the partial derivatives of the GLMM equation for $F_{n}$ and $v$ as $\frac{\partial K}{\partial F_{n}}=\alpha_{1}+\alpha_{3}v$ and $\frac{\partial K}{\partial v}=\alpha_{2}+\alpha_{3}F_{n}$.
[…]
$\%Force=\frac{C_{F}}{C_{F}+C_{V}}\times 100$ (28)
$\%Speed=\frac{C_{v}}{C_{F}+C_{V}}\times 100$ (29)
The percentage contributions of each fixed effect on the GLMM outputs for the preferred applied force $F_{n}=0.4$ N and scanning speed $v=80$ mms/s [37] for $\alpha_{1}=-0.05$, $\alpha_{2}=-4.5\times 10^{-5}$, and $\alpha_{3}=1.2\times 10^{-4}$ are then $\%Force=99.1$ and $\%Speed=0.9$.
Appendix F Spearman’s correlation between finger-electrovibration response and skin mechanics
Figure 15: Scatter plot showing the Spearman’s correlation between the finger-display interaction parameters—$K$ and $\omega_{o}$—and finger mechanical response parameters—$m$, $b$, $k$. (a) Gain, $K$ vs. finger moving mass, $m$, (b) cutoff frequency, $\omega_{o}$, vs. finger moving mass, $m$, (c) gain, $K$, vs. finger damping, $b$, (d) cutoff frequency, $\omega_{o}$, vs. finger damping, $b$, (e) gain, $K$, vs. finger stiffness, $k$, (f) cutoff frequency, $\omega_{o}$, vs. finger stiffness, $k$. The blue dots indicate the data points, and the dashed line indicates Pearson’s correlation fit. The correlation coefficient, $r$, and significance value, $p$, are shown in the plot.